40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 108     Entry time: Thu Apr 11 15:10:22 2013
Author: Koji 
Type: General 
Category: General 
Subject: OMC Progress 

[Zach, Jeff, Koji]


- Jeff configured the bottom side template to have a nominal value
obtained from the solid works model. Note that the thickness of the
curved mirrors are 6mm in the model. He added 0.3mmx2 to the dimensions.

- Jeff located the template on the breadboard such that each side has
the same amount of hanging out.

- Micrometer values

  • The one closest to the input mirror (CM1) 0.07
  • The other one on CM1 0.24
  • The one closest to the output mirror (CM2) 0.17
  • The other one on CM2 0.30
/------------\

0.17         0.07
\------------/
0.30         0.24

- Now the template is ready to accept the OMC optics.

 


- Zach and Koji finished a series of measurements for the test OMC.

Modulation depth:

- We scanned the laser PZT and recorded the data.
CH1: Reflection DC
CH2: PDH Error
CH3: Transmission (Magnified)
CH4: Transmission

- We should be able to obtain the estimation of the modulation depth and the finesse from this measurement.

- Rough calculation of the modulation depth is 0.19

Transmission:

- Incident 16.3mW
- Transmission 15.1mW
- This gave us the raw transmission of 92.6%ish.
- The modulation depth of 0.19 corresponds to 1.8% of the incident power
- The carrier reflection is almost dominated by the mode mismatch. (Note: We did not have a good resolution for the refl beam)  =>3.2%

- In total:The incident useful carrier power was 15.4mW ==> Throughput 98%
- There is slight headroom to increase the transmission by cleaning the mirrors.

FSR/Finesse:

- As our AOM is not functioning now, phase modulation sidebands are injected with the BBEOM.
- In principle, we can't expect any signal at the transmission at around the FSR frequency.
- If we apply small locking offset, the split peaks appear at the FSR frequency. The frequency of the dip corresponds to the FSR.
- We probably can extract the finesse of the cavity from this measurement. Lisa is working on this.

HOM/Finesse:

- The same PM injection gives us the frequency of the HOMs.
- We found that our EOM can work until ~500MHz.
- We could characterize the cavity resonance structure more than a single FSR.

ELOG V3.1.3-